OpenStudy (anonymous):

Write x2 + 4x = 5 in vertex form. Write your answer in the form a(x-h)^2+k=0 with no spaces

OpenStudy (solomonzelman):

\(\large\color{blue}{x^2 + 4x = 5 }\) \(\large\color{blue}{x^2 + 4x \color{red}{+4}= 5 \color{red}{+4} }\) Do you see a perfect square trinomial on the left hand side?

OpenStudy (solomonzelman):

essay?

OpenStudy (anonymous):

too bad your math teacher does not understand the difference between \[y=x^2+4x-5\] a parabola with a vertex and \[x^2+4x=5\] which is not a parabola, just an equation for which \[x=-5, x =1\]

OpenStudy (solomonzelman):

I was thinking of that too...

OpenStudy (solomonzelman):

This is not the first case, when the questions that schools give need revision.

OpenStudy (anonymous):

there is no end to bad math education not really surprising since most secondary math teacher i have known thought math was something of a mystery

OpenStudy (solomonzelman):

Anyways, As you can see, I added, 4 to both sides, \(\large\color{blue}{x^2 + 4x +4= 5 +4 }\) can you re-write the left side? (it is a perfect square trinomial)

OpenStudy (anonymous):

write them and tell them if \[x^2+4x=5\] then \(x=5\) or \(x=-1\) and that this has no vertex because it is not the graph of \[y=x^2+4x-5\]

OpenStudy (anonymous):

don't let them bulldoze you in to writing an incorrect answer just to please them

OpenStudy (solomonzelman):

well, it kind of is, it is basically a (full) function, set =0 to solve for y intercept. But, I don't know WHY they do this, either though...

OpenStudy (solomonzelman):

@Sydniariana1 Do you know what a perfect quare trinomial is?

OpenStudy (solomonzelman):

*square trinomial

OpenStudy (anonymous):

yes

OpenStudy (solomonzelman):

can you re-write the left side?

OpenStudy (solomonzelman):

\(\large\color{blue}{x^2+4x+4 }\) is \(\large\color{blue}{x^2+2x+2x+4 }\)

OpenStudy (solomonzelman):

factor that